### v1

@@ 6 Bichromatic geodesic disks
Following the procedure in [24], we obtain a similar result to Theorem 1, and we prove that it is possible to find a bichromatic pair such that any geodesic disk containing them contains at least $\left\lceil\frac{|S|-2}{72}\right\rceil$ points of $S$. Moreover, we also prove some bounds on the number of points both inside and outside the disk in this setting. These bounds have not been shown in the Euclidean setting, however, as will be shown, their proofs are similar to the proof of Theorem 5.
@@
For any set $S$ of $\frac{n}{2}\geq 3$ red points and $\frac{n}{2}\geq 3$ blue points in a polygon $\Pi^{bichrom}(n)\geq\left\lceil\frac{n-2}{72}\right\rceil\approx 0.0139n$.
@@
Let $G(S)=(V(G(S)),E(G(S)))$ be the intersection graph defined as follows: $V(S)=\{g(u,v)|u,v\in S,u\text{ is red, and }v\text{ is blue}\}$, and two vertices $g(u,v),g(x,y)$ are joined by an edge in $G(S)$ if $g(u,v)\cap g(x,y)\neq\emptyset$. The oriented graph $\overrightarrow{G}(S)$ of $G(S)$ is defined orienting an edge $g(u,v)-g(x,y)$ as $g(u,v)\rightarrow g(x,y)$ if any geodesic disk through $u$ and $v$ contains $x$ or $y$, or otherwise orienting $g(x,y)\rightarrow g(u,v)$. By Lemma 2, if two of these paths intersect, then one of the two bichromatic pairs fulfills the property that any geodesic disk through it will contain one of the other two endpoints, so this orientation is consistent.
@@
By definition of the intersection graph $G(S)$, $|E(G(S))|=I(S)$ and $|V(G(S))|=\left(\frac{n}{2}\right)^{2}$. Hence, there exists one vertex $g(u_{0},v_{0})\in\overrightarrow{G}$ with out-degree at least $\frac{I(S)}{\left(\frac{n}{2}\right)^{2}}$.
Then, we have that $\frac{I(S)}{\left(\frac{n}{2}\right)^{2}}\geq\left(\frac{\frac{n}{2}-1}{6}\right)^{2}$. This means that there exist two points $u_{0}$ and $v_{0}$, one of each color, such that the geodesic path between them intersects at least $\left(\frac{\frac{n}{2}-1}{6}\right)^{2}$ other geodesic paths. Since every point is the endpoint of at most $\frac{n}{2}-1$ geodesic paths, every geodesic disk through this pair of points contains at least $\left(\frac{\frac{n}{2}-1}{6}\right)^{2}\frac{1}{\frac{n}{2}-1}=\frac{\frac{n}{2}-1}{36}=\frac{n-2}{72}$ points of $S$. ∎
It was proved in [21] that any set of bichromatic points $S$, contains a pair $\{p,q\}\subset S$, one point of each color, such that any disk through $p$ and $q$ contains a positive fraction of the points of $S$, in particular $\frac{1}{36}$. We present the analogous colored result in the geodesic setting.
@@
For any set $S$ of $\frac{n}{2}\geq 3$ blue points and $\frac{n}{2}\geq 3$ red points in a polygon, $\Pi^{bichrom}(n)\geq\left\lceil\frac{n-2}{36}\right\rceil+2\approx 0.0277n$.
@@
$|\mathcal{Z}|\geq\frac{\dbinom{\frac{n}{2}}{3}^{2}}{\dbinom{\frac{n}{2}}{1}^{2}}=\frac{\left(\frac{n}{2}-2\right)^{2}\left(\frac{n}{2}-1\right)^{2}}{6^{2}}.$ (3)
Now we will upper-bound the value $|\mathcal{Z}|$. Let $D$ be a geodesic disk with $u$ and $v$ on its boundary, and $m-2$ points in its interior. Each $Z\in\mathcal{Z}$ contains a point of $D\setminus\{u,v\}$. The set $\{u,v\}$ can be extended to $Z$, given that $D\cap(Z\setminus\{u,v\})\neq\emptyset$, by (i) choosing a point of $Z$ from the remaining $m-2$ points in $D$, (ii) choosing a point of the same color class from the remaining $\frac{n}{2}-2$ points in $S$, and (iii) choosing two points in $S$ from the remaining color class. Thus, we get that
$|\mathcal{Z}|\leq\dbinom{m-2}{1}\dbinom{\frac{n}{2}-2}{1}\dbinom{\frac{n}{2}-1}{2}=(m-2)\frac{\left(\frac{n}{2}-2\right)^{2}\left(\frac{n}{2}-1\right)}{2}.$ (4)
@@
For any set $S$ of $\frac{n}{2}$ blue points and $\frac{n}{2}$ red points in a polygon, $\Pi^{bichrom}(n)\geq\left\lceil\frac{n}{6+\sqrt{26}}\right\rceil+1\approx\left\lceil\frac{n}{11.1}\right\rceil\approx 0.09n$.
@@
The proof is very similar to the one of Theorems 3 and 5. The difference is that we now only consider bisectors for points $p$ and $q$ of different color. The edges of the order-$1$ through order-$k$ geodesic Voronoi diagrams on $b(p,q)$ form a number of connected components (which is strictly less than the total number of edges of all the diagrams). Let $\lambda_{k}$ be the total number of connected components on all the bisectors. The goal is to find the largest value of $k$ such that $\lambda_{i}<\left(\frac{n}{2}\right)^{2}$ for all $i\in\{1,\ldots,k\}$. Since every bisector must be covered by one connected component, when $\lambda_{k}<\left(\frac{n}{2}\right)^{2}$, we note that there exists a bisector $b(p,q)$ that does not have any edges coming from the order-$1$ through order-$k$ geodesic Voronoi diagrams. Therefore, every geodesic disk with $p$ and $q$ on its boundary contains at least $k$ points.
@@
For any set $S$ of $\frac{n}{2}$ red and $\frac{n}{2}$ blue points in a polygon, $\Pi^{bichrom\mbox{-}in\mbox{-}out}(n)\geq\left\lceil\frac{n}{2(3+\sqrt{7})}\right\rceil+1\approx\left\lceil\frac{n}{11.3}\right\rceil\approx 0.0886n$.
@@
Following the proof of Theorem 5, we consider edges of both the order-$1$ through order-$k$ geodesic Voronoi diagrams, and the edges of the order-$(n-k)$ through order-$(n-1)$ geodesic Voronoi diagrams. Recall the upper bound of $\lambda_{k}+\overline{\lambda}_{n-k}\leq 3kn-2k^{2}-2k$ on the total number of connected components formed by these edges. Then as in the proof of Theorem 10, we consider only bisectors of points $p$ and $q$ of different color. There are $(\frac{n}{2})^{2}$ such bisectors, and if $\lambda_{k}+\overline{\lambda}_{n-k}<(\frac{n}{2})^{2}$, then there exists at least one bisector which does not contain an edge of the order-$1$ through order-$k$ geodesic Voronoi diagrams, or an edge of the order-$(n-k)$ through order-$(n-1)$ geodesic Voronoi diagrams. This holds as long as $k<\frac{3-\sqrt{7}}{4}n$, and the theorem follows. ∎
@@ 6.1 Upper bound
For any set $S$ of $\frac{n}{2}$ blue points and $\frac{n}{2}$ red points in a polygon, $\Pi^{bichrom}(n)\leq\left\lceil\frac{n}{5}\right\rceil+1$.
@@
We show how to construct a configuration of $n$ bichromatic points $S$ so that for every pair of bichromatic points $u,v\in S$, $\Pi^{bichrom}(n)\leq\left\lceil\frac{n}{5}\right\rceil+1$. Draw an equilateral triangle with sides of unit length and vertices labeled $x,y$, and $z$. Let $s_{1},s_{2}$ and $s_{3}$ be, respectively, the midpoints of edges ${yz},{xz}$ and ${xy}$. Let $v$ be a point on line segment ${ys_{2}}$ one unit from $y$ and farther from $s_{2}$ than from $y$, and let $w$ be a point on line segment ${xs_{1}}$ one unit from $x$ and farther from $s_{1}$ than from $x$. Place $\left\lfloor\frac{n}{5}\right\rfloor$ red points on line segment ${xy}$, and $\left\lfloor\frac{n}{5}\right\rfloor$ blue points on line segment ${zx}$. Place $\left\lfloor\frac{n}{5}\right\rfloor$ red points on line segment ${ws_{2}}$ near $w$, and $\left\lfloor\frac{n}{5}\right\rfloor$ blue points on line segment ${vs_{3}}$ near $v$ (see Figure 7). Place the remaining $\left\lfloor\frac{n}{5}\right\rfloor$ points on line segment ${zy}$. We leave it to the reader to show that through any pair of bichromatic points there is a circle containing $\left\lceil\frac{n}{5}\right\rceil+1$ points. The points can be placed in a large enough polygon without affecting the construction.
@@ 7 Conclusion
We study different variants of the following question: Given a set $S$ of $n$ points in a polygon $P$, does there always exist two points $x,y\in S$, such that every geodesic disk containing $x$ and $y$ contains a constant fraction of the points of $S$? This question has been studied intensely in the Euclidean setting  [2, 11, 12, 13, 14, 18, 21, 22, 24]. We focus on the geodesic versions of this problem. An obvious open problem is the following: Can our lower bounds be improved? Basically, some of our geodesic lower bounds are not as strong as in the Euclidean case. It would be interesting to determine whether there is a separation between the bounds in the Euclidean and geodesic case. Finally, the main open problem which has been open since 1988 is whether we can find a tight bound for the original question [18].

### v2

@@ 6 Bichromatic geodesic disks
6.1 A Constant Fraction of Bichromatic Points Inside Geodesic Disks
Following the procedure in [24], we obtain a similar result to Theorem 1, and we prove that it is possible to find a bichromatic pair such that any geodesic disk containing them contains at least $\left\lceil\frac{{n}-2}{72}\right\rceil$ points of $S$.
@@ 6.1 A Constant Fraction of Bichromatic Points Inside Geodesic Disks
For any $n\geq 6$, $\Pi^{bichrom}(n)\geq\left\lceil\frac{n-2}{72}\right\rceil\approx 0.0139n$.
@@
Analogously to Section 4.1, we define the intersection graph for the bichromatic setting as follows. Let $G(S)=(V(G(S)),E(G(S)))$ be the intersection graph defined as follows: $V(S)=\{g(u,v)\mid u,v\in S,u\text{ is red, and }v\text{ is blue}\}$, and two vertices $g(u,v),g(x,y)$ are joined by an edge in $G(S)$ if $g(u,v)$ intersects $g(x,y)$ and $u,v,x,y$ are four distinct points. The oriented graph $\overrightarrow{G}(S)$ of $G(S)$ is defined orienting an edge $\{g(u,v),g(x,y)\}$} as $g(u,v)\rightarrow g(x,y)$ if any geodesic disk through $u$ and $v$ contains $x$ or $y$, or otherwise orienting $g(x,y)\rightarrow g(u,v)$. By Lemma 2, if two of these paths intersect, then one of the two bichromatic pairs fulfills the property that any geodesic disk through it will contain one of the other two endpoints, so this orientation is consistent.
@@
By definition of the intersection graph $G(S)$, $\lvert E(G(S))\rvert=I(S)$ and $\lvert V(G(S))\rvert=\left(\frac{n}{2}\right)^{2}$. Hence, there exists one vertex $g(u_{0},v_{0})\in\overrightarrow{G}$ with out-degree at least $\frac{I(S)}{\left(\frac{n}{2}\right)^{2}}$.
Then, we have that $\frac{I(S)}{\left(\frac{n}{2}\right)^{2}}\geq\left(\frac{\frac{n}{2}-1}{6}\right)^{2}$. This means that there exist two points $u_{0}$ and $v_{0}$, one of each color, such that the geodesic path between them intersects at least $\left(\frac{\frac{n}{2}-1}{6}\right)^{2}$ other geodesic paths. Since every point is the endpoint of at most $\frac{n}{2}-1$ geodesic paths (with the minus one accounting for $u_{0}$ or $v_{0}$), every geodesic disk through this pair of points contains at least $\left(\frac{\frac{n}{2}-1}{6}\right)^{2}\frac{1}{\frac{n}{2}-1}=\frac{\frac{n}{2}-1}{36}=\frac{n-2}{72}$ points of $S$. ∎
It was proved in [21] that any set of bichromatic points $S$, contains a pair $\{p,q\}\subset S$, one point of each color, such that any disk through $p$ and $q$ contains a positive fraction of the points of $S$, in particular $\frac{1}{36}$. Although the analogous colored result in the geodesic setting does not give the optimal bound, we present it here to reflect the geodesic translation of a classical technique.
@@
For any $n\geq 6$, $\Pi^{bichrom}(n)\geq\left\lceil\frac{n-2}{36}\right\rceil+2\approx 0.0277n$.
@@
$\lvert\mathcal{Z}\rvert\geq\frac{\dbinom{\frac{n}{2}}{3}^{2}}{\dbinom{\frac{n}{2}}{1}^{2}}=\frac{\left(\frac{n}{2}-2\right)^{2}\left(\frac{n}{2}-1\right)^{2}}{6^{2}}.$ (3)
Now we will upper-bound the value $\lvert\mathcal{Z}\rvert$. Let $D$ be a geodesic disk with $u$ and $v$ on its boundary, and m2m-2 points in its interior. Each $Z\in\mathcal{Z}$ contains a point of $D\setminus\{u,v\}$. The set $\{u,v\}$ can be extended to $Z$, given that $D\cap(Z\setminus\{u,v\})\neq\emptyset$, by (i) choosing a point of $Z$ from the remaining m2m-2 points in $D$, (ii) choosing a point of the same color class from the remaining $\frac{n}{2}-2$ points in $S$, and (iii) choosing two points in $S$ from the remaining color class. Thus, we get that
$\lvert\mathcal{Z}\rvert\leq\dbinom{m-2}{1}\dbinom{\frac{n}{2}-2}{1}\dbinom{\frac{n}{2}-1}{2}=(m-2)\frac{\left(\frac{n}{2}-2\right)^{2}\left(\frac{n}{2}-1\right)}{2}.$ (4)
@@
We conclude by adapting the approach of Edelsbrunner et al. [12]. Among the classical techniques considered in this section, this method turns out to produce the strongest lower bound in the geodesic setting.
@@
For any $n\geq 12$, $\Pi^{bichrom}(n)\geq\left\lceil\frac{n}{6+\sqrt{26}}\right\rceil+1\approx\left\lceil\frac{n}{11.1}\right\rceil\approx 0.09n$.
@@
Similar to the proof of Theorems 3 and 5, we want to find a bisector that does not contain an edge of the order-$1$ through order-$k$ geodesic Voronoi diagrams, however we only consider bisectors $b(p,q)$ where $p$ and $q$ are of different color. Let $b_{k}(p,q)$ be the union of the edges of the order-$1$ through order-$k$ diagrams on $b(p,q)$. The edges of $b_{k}(p,q)$ form a number of connected components (which is strictly less than the total number of edges of all the diagrams). Let $\lambda_{k}$ be the sum of the number of connected components formed by the edges of $b_{k}(p,q)$ over all $p,q\in S$ of different color. The goal is to find the largest value of $k$ such that $\lambda_{k}<\left(\frac{n}{2}\right)^{2}$. Since every bisector must be covered by one connected component, when $\lambda_{k}<\left(\frac{n}{2}\right)^{2}$, we note that there exists a bisector $b(p,q)$ that does not have any edges coming from the order-$1$ through order-$k$ geodesic Voronoi diagrams. Therefore, every geodesic disk with $p$ and $q$ on its boundary contains at least $k$ points.
@@
6.2 A Constant Fraction of Bichromatic Points Inside and Outside Geodesic Disks
In this section, we prove a bound on the number of points both inside and outside the disk in the geodesic setting. This bound has not been shown in the Euclidean setting, however, its proof is similar to the proof of Theorem 5.
@@ 6.2 A Constant Fraction of Bichromatic Points Inside and Outside Geodesic Disks
For any $n\geq 28$, $\Pi^{bichrom\mbox{-}in\mbox{-}out}(n)\geq\left\lceil\frac{n}{14+2\sqrt{43}}\right\rceil+1\approx\left\lceil\frac{n}{27.1}\right\rceil\approx 0.0369n$.
@@
Following the proof of Theorem 5, we consider the union of the edges of both the order-$1$ through order-$k$ geodesic Voronoi diagrams, and the union of the edges of the order-(nk)(n-k) through order-(n1)(n-1) geodesic Voronoi diagrams. Recall the upper bound of $\lambda_{k}+\overline{\lambda}_{n-k}\leq 7kn-6k^{2}-6k$ on the total number of connected components formed by these edges. Then as in the proof of Theorem 10, we consider only bisectors of points $p$ and $q$ of different color. There are $(\frac{n}{2})^{2}$ such bisectors, and if $\lambda_{k}+\overline{\lambda}_{n-k}<(\frac{n}{2})^{2}$, then there exists at least one bisector which does not contain an edge of the order-$1$ through order-$k$ geodesic Voronoi diagrams, or an edge of the order-(nk)(n-k) through order-(n1)(n-1) geodesic Voronoi diagrams. This holds as long as $k<\frac{n}{14+2\sqrt{43}}$, and the theorem follows. ∎
@@ 6.3 Upper bound
For any $n\geq 6$, $\Pi^{bichrom}(n)\leq\left\lceil\frac{n}{5}\right\rceil+1$.
@@
We show how to construct a configuration of $n$ bichromatic points $S$ so that for every pair of bichromatic points $u,v\in S$, $\Pi^{bichrom}(n)\leq\left\lceil\frac{n}{5}\right\rceil+1$. Draw an equilateral triangle with sides of unit length and vertices labeled $x,y$, and $z$. Let $s_{1},s_{2}$ and $s_{3}$ be, respectively, the midpoints of edges yz,xz{yz},{xz} and xy{xy}. Let $v$ be a point on line segment $ys_{2}$} one unit from $y$ and farther from $s_{2}$ than from $y$, and let $w$ be a point on line segment $xs_{1}$ one unit from $x$ and farther from $s_{1}$ than from $x$. Place a set $S_{A}$ of $\left\lfloor\frac{n}{5}\right\rfloor$ red points on line segment xy{xy} near $y$ (distance less than $\frac{1}{4}$ is enough), and a set $S_{B}$ of $\left\lfloor\frac{n}{5}\right\rfloor$ blue points on line segment $zx$ near $x$. Place a set $S_{C}$ of $\left\lfloor\frac{n}{5}\right\rfloor$ red points on line segment $ws_{2}$ near $w$, and a set $S_{D}$ of $\left\lfloor\frac{n}{5}\right\rfloor$ blue points on line segment $vs_{3}$ near $v$. Place, in any way, a set $S_{E}$ of the remaining $\left\lfloor\frac{n}{5}\right\rfloor$ points on line segment $zy$ near $z$. The points must be placed according to this construction to achieve the desired bound. Figure 8 illustrates one possible configuration. Following the strategy in [24], we show that through any pair $p,q$ of bichromatic points there is a disk containing at most $\left\lceil\frac{n}{5}\right\rceil+1$ points. If $p,q\in S_{A}\cup S_{B}\cup S_{C}\cup S_{E}$, the result is true because of the reasoning in [24]. Moreover, if $p\in S_{D}$ we have three cases: (i) if $q\in S_{A}$, the disk through $p$ and $q$ and tangent to the line $xy$ contains at most all the points in $S_{D}$, plus the point $q$, (ii) if $q\in S_{C}$, the disk through $p$ and $q$ and tangent to the line $vs_{3}$ contains at most all the points in $S_{C}$, plus the point $q$, and (iii) if $q\in S_{E}$, the disk through $p$ and $q$ and tangent to the line $zy$ contains at most all the points in $S_{D}$, plus the point $q$. The points can be placed in a large enough polygon without affecting the construction.
@@ 7 Conclusion
We study different variants of the following question: Given a set $S$ of $n$ points in a polygon $P$, does there always exist two points $x,y\in S$, such that every geodesic disk containing $x$ and $y$ contains a constant fraction of the points of $S$? This question has been studied intensely in the Euclidean setting [2, 11, 12, 14, 13, 18, 21, 22, 24]. We focus on the geodesic versions of this problem. An obvious open problem is the following: Can our lower bounds be improved? Basically, some of our geodesic lower bounds are not as strong as in the Euclidean case. It would be interesting to determine whether there is a separation between the bounds in the Euclidean and geodesic case. Additionally, in the Euclidean setting the best-known lower bound for the inside-outside version of the problem match the best known lower bounds for the original version of the problem exactly, whereas we proved different lower bounds in this paper. Is this difference simply an artifact of our proof or do these two versions of the problem actually have different answers? Note that even in the Euclidean case it is not known whether the two versions of the problem have different answers or not. Finally, the main open problem which has been open since 1988 is whether we can find a tight bound for the original question [18].